You are given results on a case control study on alcoholism and breast cancer, 200m participants with Breast cancer and 200 participants without cancer were selected and questioned on previous alcohol use.

Breast Cancer (+)

Breast Cancer (-)

Total

Heavy alcohol use

40

20

60

No/low alcohol use

160

180

340

total

200

200

400

  1. Calculate and interpret the appropriate measure of association for this study?
  2. You later learn that 15 of the controls [Breast CANCER (-)] did not remember their previous heavy alcohol use and incorrectly reported that they were low alcohol users. By comparison all the breast cancer (+) cases correctly remembered their alcohol use. Please compare appropriate measure of association. (HINT: create a new 2x2 table below)
  3. Was there evidence of Bias in this study? If so what type of Bias?
Answer #1

Yule’s coefficient of Association ‘Q’ is given by

(AB (ab)-(Ab (aB) (AB) (ab)+(Ab)(aB) 0.384615385=Q1(say) (using table 1)

As Q is >0 implies there is a positive association between 'Breast cancer' and "use of alcohol'

Table 1

Breast Cancer (+) denoted 'A' Breast Cancer (-) denoted 'a' Total
Heavy alcohol use denoted 'B' 40 (AB) 20 (aB) 60 (B)
No/low alcohol use denoted 'b' 160 (Ab) 180 (ab) 340 (b)
total 200 (A) 200 (a) 400 (N)

As it is reported that 15 of the controls [Breast CANCER (-)] did not remember their previous heavy alcohol use and incorrectly reported that they were low alcohol users. So corrected frequency table is shown in Table 2:

(AB (ab)-(Ab (aB) (AB) (ab)+(Ab)(aB)    0.081967213 =Q2 (corrected) (using table 2)

Table 2
Breast Cancer (+) denoted 'A' Breast Cancer (-) denoted 'a' Total
Heavy alcohol use denoted 'B' 40 (AB) 35 (aB) 60 (B)
No/low alcohol use denoted 'b' 160 (Ab) 180 (ab) 340 (b)
total 200 (A) 165 (a) 400 (N)

Obviously there exists a Bias in this study as Q1 and Q2 are not same and this types of bias is called misclassification bias. Misclassification bias is a types of bias that appear due to negligence of data collection.